OpenStudy (anonymous):

Identify the vertex, axis of symmetry, maximum and minimum value, and range of y=3x^2 -6x+2 Can someone please help me find these? Thanks in advance.

OpenStudy (wolf1728):

The coefficient of x² is 3 which is positive so that means it will have a curve that is concave up (shaped like a 'U'). It will have a minimum value and its maximum value will be infinity.

OpenStudy (anonymous):

Okay, thanks. But how did you get that?

OpenStudy (wolf1728):

Just by knowing that the coefficient of x² is a positive number '3'.

OpenStudy (anonymous):

I got... vertex: x=- 3/2(-6) = 1/4 y=3(1/4)^2 -6(1/4)+2 = 11/16 axis of symmetry: x= 1/4 minimum value: y= 11/16 at x= 1/4 range: y (greater than or equal to) 11/16 Where did I go wrong?

OpenStudy (wolf1728):

It seems you'll also need the vertex which will also give you a minimum value. This is from calculus but gives you the correct answer. We take the derivative of y=3x^2 -6x+2 and make it equal zero. The derivative is 6x -6 and 6x - 6 = 0 when x = 1. The minimum value is at x=1 and when x = 1, y= 3x² -6x +2 = 3-6+2 = -1 So minimum is at (1, -1) The minimum occurs at (1, -1) so the axis of symmetry is a vertical line going through the point x = 1.

OpenStudy (anonymous):

y = (3x)squared - 4x - 2 well you can do the completing the squares so factor this to: y = ax(sqared) + bx + c ==> y = a(x-h)^2 + k y = (3x)squared - 4x +2 y = (3x^2 - 4x +4)+ 2 - 4 =============> ((1/2)(-4))) = (-2)^2 = +4 factor the inside!!! Use quadratic do the rest srry dude to lazy D: hope this helps

OpenStudy (anonymous):

@Gracy_KSL @wolf1728 thanks both of you. So what exactly did I do wrong?

OpenStudy (wolf1728):

I'd guess that you used the vertex formula incorrectly. (I don't know it - I can look it up).

OpenStudy (anonymous):

That's okay. Thanks so much!! :)

OpenStudy (wolf1728):

Okay the vertex formula for the x value is x = -b/2a In the quadratic equation 3x^2 -6x+2 a=3 b=-6 anc c=2 x = -(-6) / 2*3 x = 6/6 x=1

OpenStudy (wolf1728):

I'd guess you get the y value by plugging in the value of x =1 into the equation = 3*1 -6*1 +2 = -1

OpenStudy (anonymous):

Thank you so much. You were very helpful!!

OpenStudy (wolf1728):

Thanks Karebear :-)